### v1

@@ Abstract
The discovery and peculiarities of hot Jupiters present significant challenges to the classical planet formation theories based on our Solar system [1, 2]. To explain the origin of hot Jupiters, previous studies have proposed various models, e.g., in-situ formation [3, 4], disk migration [5, 6] and high-eccentricity migration [7, 8, 9, 10, 11]. However, none of these models alone can satisfactorily explain all the observational evidence, and the origin of hot Jupiters remains puzzling (see the review by [2]). An outstanding question is: what is the predominant channel and how much do different models contribute to the formation of hot Jupiters?
@@
Furthermore, as hot Jupiters are close to the host stars, their orbits are expected to decay with age due to tidal interaction [12, 13]. Previous studies have looked for observational evidence for the tidal decays from individual [14, 15, 16] and ensemble properties of hot Jupiters [12, 17, 18, 19, 20] as well as the properties (e.g., age, spin-up) of their host stars [19, 21, 22, 23, 24]. However, the derived stellar modified tidal quality factor $Q^{{}^{\prime}}_{*}$ from different works vary by several orders of magnitude [12, 17, 18, 19, 14, 15, 25], ranging from $\sim 10^{4}-10^{9}$ (Extended Data Tab. 1). Therefore, obtaining a well-constrained stellar tidal quality factor is vital to resolve this discrepancy and to achieve a fundamental understanding of tidal phenomena.
@@
The age-frequency relationship of hot Jupiters offers novel insights into their origin and tidal evolution. Specifically, different formation mechanisms operate on timescales ranging from millions to billions of years [2], resulting in distinct arrival time distributions. After hot Jupiters arrive at $\lesssim 0.1$ AU and have been largely circularized, the orbits of hot Jupiters are expected to decay under tidal forces, with the decay rate primarily determined by $Q^{{}^{\prime}}_{*}$ [12, 13]. Some of them will eventually pass into the Roche limit and be disrupted within the lifetime of their host stars [13, 21]. Consequently, hot Jupiters formed through different origin models, under the tidal forces of stars with varying $Q^{{}^{\prime}}_{*}$, will exhibit distinct evolutionary patterns over time.
@@
Based on the kinematic method for estimating age and a sample of hot Jupiters with kinematic properties, in a previous study [21], we derived the age-frequency relation and revealed a generally declining trend using a mathematical exponential function. In this work, we perform detailed analyses of the age-frequency trend and focus on the physical interpretation of the observational results by quantitatively fitting with the theoretical expectations of the tidal evolution of hot Jupiters originating from different channels. This not only allows us to obtain a precise constraint on the stellar tidal quality factor but also enables the determination of the relative importance of different models in forming hot Jupiters across varying timescales.
@@
Our work builds upon the planetary sample selected and the stellar parent sample constructed in our previous study [for details, see the § 1 and 3 of Supporting information of 21]. The planetary sample was collected from the kinematic catalogs of the Planets Across Space and Time series [PAST; 26, 27] and composed of giant planets orbiting single Sun-like stars in the Galactic disks, which was further divided into three subsamples based on the discovery method and facility, i.e., radial velocity (RV), space-based transit (ST), and ground-based transit (GT). In this work, we further expand the GT subsample by utilizing the Gaia DR2 catalog. The final sample consists of 123 hot Jupiters (17 RV, 14 ST and 92 GT) and 936,815 (1662 RV, 19,228 ST and 915,925 GT) parent stars (Methods, § 1.1).
@@
We derive a ‘combined’ frequency of hot Jupiters $F_{\rm HJ}$ (i.e., also termed as ’occurrence’, defined as the fraction of stars with a hot Jupiter) as a function of kinematic age by jointly using the data from the three subsamples discovered by different methods/facilities (Methods, § 1.2). The detection efficiencies and geometric effects are corrected with the same procedures described in § 3 of Supporting information of [21]. As can be seen in the top panel of Figure 1, $F_{\rm HJ}$ generally declines with increasing age (Methods, § 1.3.1), which is consistent with previous studies [21, 22].
@@
Various mechanisms have been proposed to explain the origin of hot Jupiters [2]. Considering formation timescales as well as age uncertainties ($\gtrsim$ tenth of Gyr), in this work, these mechanisms are divided into two categories: ‘Early’ model and ‘Late’ model (Methods, § 2.1). For the ‘Early’ model, most hot Jupiters form within about several tenths of Gyr via in-situ formation [3, 4], disk migration [5, 6], planet-planet scattering[7]/Kozai [9]/coplanar secular [11]. Whereas, for ‘Late’ model, hot Jupiters are delivered to $\lesssim 0.1$ AU on a long timescale up to several Gyrs. Although the planet-star Kozai mechanism can potentially deliver hot Jupiters over Gyr [28, 29], we do not consider it as the main channel for the following reasons: (1) the fraction of binary companions capable of inducing Kozai–Lidov oscillations is $\lesssim 16\%\pm 5\%$, and the upper limit will further decrease to $6\%\pm 2\%$ assuming an isotropic distribution for the stellar companion inclinations [30]; (2) the typical timescales for planet-star Kozai cycles are predominantly (99.71%) less than 1 Gyr (Extended Data Fig. 1); (3) potential binaries have been largely excluded in our sample. Instead, secular chaos can be a promising candidate for the ‘Late’ model, serving as a robust mechanism consistently capable of producing hot Jupiters around single stars over Gyr timescales. Previous numerical studies have predicted the formation rate of hot Jupiters through secular chaos as a function of stellar age [31]. We also compared the arrival timescales of the secular chaos model with the estimated arrival timescales of the observed‘late-arrived’ population, which are statistically consistent (with KS $p$-value of 0.3499) after considering the impact of selection criteria (Extended Data Fig. 2). In the following, we test the secular chaos mechanism as the ‘Late’ model.
@@
In addition, previous studies have suggested that the stellar $Q^{{}^{\prime}}_{*}$ may depend on the stellar mass and the evolutionary stage [32, 24, 33] as well as the orbital period $P$ of hot Jupiters [24]. Since we only study single Sun-like stars, these stellar properties are similar across different age bins (Figure S12). To account for this potential $Q^{{}^{\prime}}_{*}-P$ dependence, we also fit the observational data using an empirical formula and find that the hybrid model remains preferred compared to the single-‘Early’ model and single-‘Late’ model, with confidence levels >4>4-$\sigma$ (Figure S13-15). Besides, we also adopted the secular chaos model from [34] as the ‘late’ model, and the derived best-fits of $f_{\rm Late}$ and $Q^{{}^{\prime}}_{*}$ are highly comparable (Figure S16), suggesting that our results are not sensitive to the choice of secular chaos model.
@@
The agreement between the observed age-frequency relation and the hybrid model implies that secular chaos [10, 31, 34] is the dominant mechanism to form the late arrived hot Jupiter. We further explore other properties of hot Jupiter systems and find some additional evidence supporting this inference. As shown in Figure 3, we find that the ‘late-arrived’ hot Jupiter systems have a stellar obliquity distribution mainly within $10-90^{\circ}$, which is consistent with the prediction of the secular chaos mechanism [34] with a KS $p$-value of 0.229, but it deviates from the predictions of all other mechanisms with KS $p$-values<0.05<0.05 (Methods, § 3.2, Extended Data Tab. 3). In addition, since secular chaos mechanism occurs in systems with multiple giant planets, one would expect that the ‘late-arrived’ hot Jupiters are more likely to have outer companions and are preferentially hosted by metal-rich stars. For consistency check, we investigate their metallicity distribution and companion presence. As expected, the ‘late-arrived’ hot Jupiter systems are metal-richer ($\rm\overline{[Fe/H]}=0.12^{+0.02}_{-0.02}$ dex) (Methods, § 3.3, Extended Data Fig. 7) and show a larger fraction ($F_{\rm OG}=40^{+26.7}_{-23.3}\%$; Methods, § 3.4) to have outer giant companions compared to the ‘early-arrived’ hot Jupiters ($\rm\overline{[Fe/H]}=0.06^{+0.02}_{-0.03}$, $F_{\rm OG}=8.3^{+7.6}_{-7.0}\%$).
@@
To further validate the constraints on $\log Q^{{}^{\prime}}_{*}$ derived in our work, we calculate the absolute decay rates $|\dot{P}|$ of orbital period and the expected numbers of hot Jupiters with detectable orbital period decay, $N^{\rm detectable}_{\rm decay}$ (see Figure 4), and then compare these with the observation number, i.e., 1 confirmed case +10 candidates identified in Wang et al. (2024) ([16], Methods, § 3.5). As can be seen, the expected value $N^{\rm detectable}_{\rm decay}=4\pm 2$ (red lines), which lies within the observed range of 1111-11 with a confidence level of 99.5%, aligning well with the observational results. Previous studies have also provided constraints on $Q^{{}^{\prime}}_{*}$ (Extended Data Tab. 1) based on the evolution of hot Jupiters in the individual systems [14, 16] and the ensemble statistics [12, 17, 18, 19, 20]. For comparison, we tested $Q^{{}^{\prime}}_{*}$ derived from previous studies by repeating the above procedure and found that the resulted $N^{\rm detectable}_{\rm decay}$ do not match the observations. Specifically, when taking the constraints of $\log Q^{{}^{\prime}}_{*}=5.14\pm 0.05$ derived from individual system WASP 12 (the only one with conclusive evidence of orbital decay; green lines), we found that $N^{\rm detectable}_{\rm decay}=15\pm 3$, which is significantly ($\gtrsim 4$-$\sigma$) higher than the number of confirmed system and remains overestimated even if all 10 candidates are eventually confirmed. In contrast, using the constraints of $\log Q^{{}^{\prime}}_{*}\sim 6-9$ derived from previous ensemble studies (blue lines), the orbital decay of hot Jupiters is expected to be unobservable yet, with a confidence level of 92.5%.
@@
One of the longest-standing puzzles in exoplanet research is how hot Jupiters form. Previous studies have attempted to constrain the origin of hot Jupiters from the observed distribution of the semi-major axis/eccentricity/age/configurations for hot Jupiters [see the review by 2]. For example, Wu et al. (2023) [35] reporetd that $\gtrsim 12\pm 6\%$ of hot Jupiters have nearby planetary companions and thus form via quiescent dynamical mechanisms (i.e., in-situ formation and disk migration). Hamer et al. (2022) [36] found that high-stellar-obliquity hot Jupiter systems are kinematically younger than low-stellar-obliquity hot Jupiter systems, implying that some hot Jupiters in misaligned systems may arrive late. However, it remains unclear which of these formation mechanisms actually plays a role and how much they quantitatively contribute [2]. In this study, we analyze the age-frequency relation of hot Jupiters orbiting single Sun-like stars and provide direct evidence that hot Jupiters form through multiple channels operating on different timescales (Figure 1). We further quantitatively determine the contributions of different channels, revealing that $38^{+16}_{-14}\%$ of hot Jupiters should arrive late with timescales up to several Gyrs via the ‘Late’ model (Figure 2). Additionally, we found more evidence from obliquity that supports secular chaos [10, 31, 34] as a promising mechanism for producing such a ‘late-arrived’ population (Figure 3).
@@
The evolutionary patterns of hot Jupiters have long served as a critical framework for testing tidal theories and constraining stellar tidal factors (Extended Data Tab. 1). However, $Q^{{}^{\prime}}_{*}$ remain poorly constrained, with notable discrepancies between results obtained from the orbital decay of hot Jupiter/spin-up of host stars in individual systems $\sim 4.5-7$; 14, 16, 24] and the observed distributions of hot Jupiter ensembles $\sim 6-9$; 12, 17, 18, 19, 20]. Moreover, the $Q^{{}^{\prime}}_{*}$ values from previous studies do not match the observed number of hot Jupiters with orbital decay (Figure 4). This is likely due to the limitations of their methodologies. Specifically, the limited sample size of individual systems with precise orbital decay rates or stellar spin periods/ages hampers the ability to represent the overall distribution, potentially introducing a bias toward a larger $|\dot{P}|$ and lower tail of $Q^{{}^{\prime}}_{*}$. On the other hand, previous ensemble studies on the distribution/evolution of the hot Jupiter population [12, 17, 18, 19, 20] assumed that hot Jupiters were all ‘early-arrived’ and neglected the contribution of ‘late-arrived’ population, leading to overestimate in $Q^{{}^{\prime}}_{*}$ and thus an underestimate in $N^{\rm detectable}_{\rm decay}$. Hansen 2010 [37] and 2012 [38] constrained $\log Q^{{}^{\prime}}_{*}\sim 7-9$ from the eccentricity distribution of short-period exoplanets, but the analysis involved host stars near the convective-radiative envelope transition that may have less efficient dissipation and larger $Q^{{}^{\prime}}_{*}$ [24]. In this work, by considering the effects of not only early formed but also late-arrived hot Jupiters, we obtain a more precise constraint on $\log Q^{{}^{\prime}}_{*}=5.7^{+0.4}_{-0.3}$ for Sun-like stars from a large sample analysis, aligning well with both the observed age-frequency relation (Figure 2) and the detected number of hot Jupiters with orbital decay (red lines in Figure 4). Very recently, during the peer-review period of this paper, we noted that Millholland et al. (2025) [39] analyzed the distribution of insprial timescales ($\sim 0.001-0.3$ Gyr) of short-period hot Jupiters and provided a constraint of $\log Q^{*}_{{}^{\prime}}\sim 5.5-7.0$, consistent with our results. Such a consistence is expected since these hot Jupiters are probably in the very late stages of tidal evolution, which is primarily dominated by $Q^{{}^{\prime}}_{*}$ rather than formation channels. Furthermore, our derived constraint is well consistent with those derived from binary interactions ($\log Q^{{}^{\prime}}_{*}\sim 5.5$; [40, 41]). In addition, our results emphasizes the necessity of considering the formation pathways and timescales of hot Jupiters when constraining $Q^{{}^{\prime}}_{*}$.
@@
In summary, we find that the declining of the frequency of hot Jupiters with age is broken, which simultaneously puts constraints on the stellar tidal quality factors and the relative importance of the ‘Early’ and ‘Late’ models to form hot Jupiters. Although the secular chaos mechanism [10, 31, 34] is identified as the dominant channel in the ‘Late’ model, the situation regarding the ‘Early’ model (timescales within tenths of Gyr) remains unclear, highlighting the need for a larger sample of hot Jupiters orbiting young stars. Moreover, our results focus primarily on understanding formation channels in single-star systems. Investigating planetary systems in binary star samples can offer further insights into constraining the formation mechanisms (e.g., binary Kozai) of hot Jupiters [e.g., 42, 43]. Future observations and analyses of hot Jupiters from TESS [44], PLATO [45], JWST [46] and ET 2.0 [47] will test our results and further provide more insights on their origin and evolution.

### v2

@@ Abstract
The discovery and peculiarities of hot Jupiters present significant challenges to the classical planet formation theories based on our Solar system [1995Natur.378..355M, 2018ARA&A..56..175D]. To explain the origin of hot Jupiters, previous studies have proposed various models, e.g., in-situ formation [2016ApJ...829..114B, 2016ApJ...817L..17B], disk migration [1996Natur.380..606L, 2014prpl.conf..667B] and high-eccentricity migration [1996Sci...274..954R, 2003ApJ...589..605W, 2011Natur.473..187N, 2011ApJ...735..109W, 2015ApJ...805...75P]. However, none of these models alone can satisfactorily explain all the observational evidence, and the origin of hot Jupiters remains puzzling (see the review by [2018ARA&A..56..175D]). An outstanding question is: what is the predominant channel and how much do different models contribute to the formation of hot Jupiters?
@@
Furthermore, as hot Jupiters are close to the host stars, their orbits are expected to decay with age due to tidal interaction [2008ApJ...678.1396J, 2012MNRAS.423..486L]. Previous studies have looked for observational evidence for the tidal decays from individual [2020ApJ... [^8] L...5Y, 2021AJ....161...72T, 2024ApJS..270...14W] and ensemble properties of hot Jupiters [2008ApJ...678.1396J, 2017AA...602A.107B, 2018MNRAS.476.2542C, 2019AJ....158..190H, 2020AJ....160..138H] as well as the properties (e.g., age, spin-up) of their host stars [2019AJ....158..190H, 2023PNAS..12004179C, 2023AJ....166..209M, 2024AJ....168....7B, 2018AJ....155..165P]. However, the derived stellar modified tidal quality factor $Q^{{}^{\prime}}_{*}$ from different works vary by several orders of magnitude [2008ApJ...678.1396J, 2017AA...602A.107B, 2018MNRAS.476.2542C, 2019AJ....158..190H, 2020ApJ... [^8] L...5Y, 2021AJ....161...72T, 2022AJ....163..208B], ranging from $\sim 10^{4}-10^{9}$ (Extended Data Tab. 1). Therefore, obtaining a well-constrained stellar tidal quality factor is vital to resolve this discrepancy and to achieve a fundamental understanding of tidal phenomena.
@@
The age-frequency relationship of hot Jupiters offers novel insights into their origin and tidal evolution. Specifically, different formation mechanisms operate on timescales ranging from millions to billions of years [2018ARA&A..56..175D], resulting in distinct arrival time distributions. After hot Jupiters arrive at $\lesssim 0.1$ AU and have been largely circularized, the orbits of hot Jupiters are expected to decay under tidal forces, with the decay rate primarily determined by $Q^{{}^{\prime}}_{*}$ [2008ApJ...678.1396J, 2012MNRAS.423..486L]. Some of them will eventually pass into the Roche limit and be disrupted within the lifetime of their host stars [2012MNRAS.423..486L, 2023PNAS..12004179C]. Consequently, hot Jupiters formed through different origin models, under the tidal forces of stars with varying $Q^{{}^{\prime}}_{*}$, will exhibit distinct evolutionary patterns over time.
@@
Based on the kinematic method for estimating age and a sample of hot Jupiters with kinematic properties, in a previous study [2023PNAS..12004179C], we derived the age-frequency relation and revealed a generally declining trend using a mathematical exponential function. In this work, we perform detailed analyses of the age-frequency trend and focus on the physical interpretation of the observational results by quantitatively fitting with the theoretical expectations of the tidal evolution of hot Jupiters originating from different channels. This not only allows us to obtain a precise constraint on the stellar tidal quality factor but also enables the determination of the relative importance of different models in forming hot Jupiters across varying timescales.
@@
Our work builds upon the planetary sample selected and the stellar parent sample constructed in our previous study [for details, see the § 1 and 3 of Supporting information of 2023PNAS..12004179C]. The planetary sample was collected from the kinematic catalogs of the Planets Across Space and Time series [PAST; 2021ApJ...909..115C, 2021AJ....162..100C] and composed of giant planets orbiting single Sun-like stars in the Galactic disks, which was further divided into three subsamples based on the discovery method and facility, i.e., radial velocity (RV), space-based transit (ST), and ground-based transit (GT). In this work, we further expand the GT subsample by utilizing the Gaia DR2 catalog. The final sample consists of 123 hot Jupiters (17 RV, 14 ST and 92 GT) and 936,815 (1662 RV, 19,228 ST and 915,925 GT) parent stars (Methods, § 1.1).
@@
We derive a ‘combined’ frequency of hot Jupiters $F_{\rm HJ}$ (i.e., also termed as ’occurrence’, defined as the fraction of stars with a hot Jupiter) as a function of kinematic age by jointly using the data from the three subsamples discovered by different methods/facilities (Methods, § 1.2). The detection efficiencies and geometric effects are corrected with the same procedures described in § 3 of Supporting information of [2023PNAS..12004179C]. As can be seen in the top panel of Figure 1, $F_{\rm HJ}$ generally declines with increasing age (Methods, § 1.3.1), which is consistent with previous studies [2023PNAS..12004179C, 2023AJ....166..209M].
@@
Various mechanisms have been proposed to explain the origin of hot Jupiters [2018ARA&A..56..175D]. Considering formation timescales as well as age uncertainties ($\gtrsim$ tenth of Gyr), in this work, these mechanisms are divided into two categories: ‘Early’ model and ‘Late’ model (Methods, § 2.1). For the ‘Early’ model, most hot Jupiters form within about several tenths of Gyr via in-situ formation [2016ApJ...829..114B, 2016ApJ...817L..17B], disk migration [1996Natur.380..606L, 2014prpl.conf..667B], planet-planet scattering[1996Sci...274..954R]/Kozai [2011Natur.473..187N]/coplanar secular [2015ApJ...805...75P]. Whereas, for ‘Late’ model, hot Jupiters are delivered to $\lesssim 0.1$ AU on a long timescale up to several Gyrs. Although the planet-star Kozai mechanism can potentially deliver hot Jupiters over Gyr [2003ApJ...589..605W, 2007ApJ...669.1298F], we do not consider it as the main channel for the following reasons: (1) the fraction of binary companions capable of inducing Kozai–Lidov oscillations is $\lesssim 16\%\pm 5\%$, and the upper limit will further decrease to $6\%\pm 2\%$ assuming an isotropic distribution for the stellar companion inclinations [2016ApJ...827....8N]; (2) the typical timescales for planet-star Kozai cycles are predominantly (99.71%) less than 1 Gyr (Extended Data Fig. 1); (3) potential binaries have been largely excluded in our sample. Instead, secular chaos can be a promising candidate for the ‘Late’ model, serving as a robust mechanism consistently capable of producing hot Jupiters around single stars over Gyr timescales. Previous numerical studies have predicted the formation rate of hot Jupiters through secular chaos as a function of stellar age [2017MNRAS.464..688H]. We also compared the arrival timescales of the secular chaos model with the estimated arrival timescales of the observed‘late-arrived’ population, which are statistically consistent (with KS $p$-value of 0.3499) after considering the impact of selection criteria (Extended Data Fig. 2). In the following, we test the secular chaos mechanism as the ‘Late’ model.
@@
In addition, previous studies have suggested that the stellar $Q^{{}^{\prime}}_{*}$ may depend on the stellar mass and the evolutionary stage [2017ApJ...849L..11W, 2018AJ....155..165P, 2020MNRAS.498.2270B] as well as the orbital period $P$ of hot Jupiters [2018AJ....155..165P]. Since we only study single Sun-like stars, these stellar properties are similar across different age bins (Figure S12). To account for this potential $Q^{{}^{\prime}}_{*}-P$ dependence, we also fit the observational data using an empirical formula and find that the hybrid model remains preferred compared to the single-‘Early’ model and single-‘Late’ model, with confidence levels >4>4-$\sigma$ (Figure S13-15). Besides, we also adopted the secular chaos model from [2019MNRAS.486.2265T] as the ‘late’ model, and the derived best-fits of $f_{\rm Late}$ and $Q^{{}^{\prime}}_{*}$ are highly comparable (Figure S16), suggesting that our results are not sensitive to the choice of secular chaos model.
@@
The agreement between the observed age-frequency relation and the hybrid model implies that secular chaos [2011ApJ...735..109W, 2017MNRAS.464..688H, 2019MNRAS.486.2265T] is the dominant mechanism to form the late arrived hot Jupiter. We further explore other properties of hot Jupiter systems and find some additional evidence supporting this inference. As shown in Figure 3, we find that the ‘late-arrived’ hot Jupiter systems have a stellar obliquity distribution mainly within $10-90^{\circ}$, which is consistent with the prediction of the secular chaos mechanism [2019MNRAS.486.2265T] with a KS $p$-value of 0.229, but it deviates from the predictions of all other mechanisms with KS $p$-values<0.05<0.05 (Methods, § 3.2, Extended Data Tab. 3). In addition, since secular chaos mechanism occurs in systems with multiple giant planets, one would expect that the ‘late-arrived’ hot Jupiters are more likely to have outer companions and are preferentially hosted by metal-rich stars. For consistency check, we investigate their metallicity distribution and companion presence. As expected, the ‘late-arrived’ hot Jupiter systems are metal-richer ($\rm\overline{[Fe/H]}=0.12^{+0.02}_{-0.02}$ dex) (Methods, § 3.3, Extended Data Fig. 7) and show a larger fraction ($F_{\rm OG}=40^{+26.7}_{-23.3}\%$; Methods, § 3.4) to have outer giant companions compared to the ‘early-arrived’ hot Jupiters ($\rm\overline{[Fe/H]}=0.06^{+0.02}_{-0.03}$, $F_{\rm OG}=8.3^{+7.6}_{-7.0}\%$).
@@
To further validate the constraints on $\log Q^{{}^{\prime}}_{*}$ derived in our work, we calculate the absolute decay rates $|\dot{P}|$ of orbital period and the expected numbers of hot Jupiters with detectable orbital period decay, $N^{\rm detectable}_{\rm decay}$ (see Figure 4), and then compare these with the observation number, i.e., 1 confirmed case +10 candidates identified in Wang et al. (2024) ([2024ApJS..270...14W], Methods, § 3.5). As can be seen, the expected value $N^{\rm detectable}_{\rm decay}=4\pm 2$ (red lines), which lies within the observed range of 1111-11 with a confidence level of 99.5%, aligning well with the observational results. Previous studies have also provided constraints on $Q^{{}^{\prime}}_{*}$ (Extended Data Tab. 1) based on the evolution of hot Jupiters in the individual systems [2020ApJ... [^8] L...5Y, 2024ApJS..270...14W] and the ensemble statistics [2008ApJ...678.1396J, 2017AA...602A.107B, 2018MNRAS.476.2542C, 2019AJ....158..190H, 2020AJ....160..138H]. For comparison, we tested $Q^{{}^{\prime}}_{*}$ derived from previous studies by repeating the above procedure and found that the resulted $N^{\rm detectable}_{\rm decay}$ do not match the observations. Specifically, when taking the constraints of $\log Q^{{}^{\prime}}_{*}=5.14\pm 0.05$ derived from individual system WASP 12 (the only one with conclusive evidence of orbital decay; green lines), we found that $N^{\rm detectable}_{\rm decay}=15\pm 3$, which is significantly ($\gtrsim 4$-$\sigma$) higher than the number of confirmed system and remains overestimated even if all 10 candidates are eventually confirmed. In contrast, using the constraints of $\log Q^{{}^{\prime}}_{*}\sim 6-9$ derived from previous ensemble studies (blue lines), the orbital decay of hot Jupiters is expected to be unobservable yet, with a confidence level of 92.5%.
@@
One of the longest-standing puzzles in exoplanet research is how hot Jupiters form. Previous studies have attempted to constrain the origin of hot Jupiters from the observed distribution of the semi-major axis/eccentricity/age/configurations for hot Jupiters [see the review by 2018ARA&A..56..175D]. For example, Wu et al. (2023) [2023AJ....165..171W] reporetd that $\gtrsim 12\pm 6\%$ of hot Jupiters have nearby planetary companions and thus form via quiescent dynamical mechanisms (i.e., in-situ formation and disk migration). Hamer et al. (2022) [2022AJ....164...26H] found that high-stellar-obliquity hot Jupiter systems are kinematically younger than low-stellar-obliquity hot Jupiter systems, implying that some hot Jupiters in misaligned systems may arrive late. However, it remains unclear which of these formation mechanisms actually plays a role and how much they quantitatively contribute [2018ARA&A..56..175D]. In this study, we analyze the age-frequency relation of hot Jupiters orbiting single Sun-like stars and provide direct evidence that hot Jupiters form through multiple channels operating on different timescales (Figure 1). We further quantitatively determine the contributions of different channels, revealing that $38^{+16}_{-14}\%$ of hot Jupiters should arrive late with timescales up to several Gyrs via the ‘Late’ model (Figure 2). Additionally, we found more evidence from obliquity that supports secular chaos [2011ApJ...735..109W, 2017MNRAS.464..688H, 2019MNRAS.486.2265T] as a promising mechanism for producing such a ‘late-arrived’ population (Figure 3).
@@
The evolutionary patterns of hot Jupiters have long served as a critical framework for testing tidal theories and constraining stellar tidal factors (Extended Data Tab. 1). However, $Q^{{}^{\prime}}_{*}$ remain poorly constrained, with notable discrepancies between results obtained from the orbital decay of hot Jupiter/spin-up of host stars in individual systems $\sim 4.5-7$; 2020ApJ... [^8] L...5Y, 2024ApJS..270...14W, 2018AJ....155..165P] and the observed distributions of hot Jupiter ensembles $\sim 6-9$; 2008ApJ...678.1396J, 2017AA...602A.107B, 2018MNRAS.476.2542C, 2019AJ....158..190H, 2020AJ....160..138H]. Moreover, the $Q^{{}^{\prime}}_{*}$ values from previous studies do not match the observed number of hot Jupiters with orbital decay (Figure 4). This is likely due to the limitations of their methodologies. Specifically, the limited sample size of individual systems with precise orbital decay rates or stellar spin periods/ages hampers the ability to represent the overall distribution, potentially introducing a bias toward a larger $|\dot{P}|$ and lower tail of $Q^{{}^{\prime}}_{*}$. On the other hand, previous ensemble studies on the distribution/evolution of the hot Jupiter population [2008ApJ...678.1396J, 2017AA...602A.107B, 2018MNRAS.476.2542C, 2019AJ....158..190H, 2020AJ....160..138H] assumed that hot Jupiters were all ‘early-arrived’ and neglected the contribution of ‘late-arrived’ population, leading to overestimate in $Q^{{}^{\prime}}_{*}$ and thus an underestimate in $N^{\rm detectable}_{\rm decay}$. Hansen 2010 [2010ApJ...723..285H] and 2012 [2012ApJ...757....6H] constrained $\log Q^{{}^{\prime}}_{*}\sim 7-9$ from the eccentricity distribution of short-period exoplanets, but the analysis involved host stars near the convective-radiative envelope transition that may have less efficient dissipation and larger $Q^{{}^{\prime}}_{*}$ [2018AJ....155..165P]. In this work, by considering the effects of not only early formed but also late-arrived hot Jupiters, we obtain a more precise constraint on $\log Q^{{}^{\prime}}_{*}=5.7^{+0.4}_{-0.3}$ for Sun-like stars from a large sample analysis, aligning well with both the observed age-frequency relation (Figure 2) and the detected number of hot Jupiters with orbital decay (red lines in Figure 4). Very recently, during the peer-review period of this paper, we noted that Millholland et al. (2025) [2025ApJ...981...77M] analyzed the distribution of insprial timescales ($\sim 0.001-0.3$ Gyr) of short-period hot Jupiters and provided a constraint of $\log Q^{*}_{{}^{\prime}}\sim 5.5-7.0$, consistent with our results. Such a consistence is expected since these hot Jupiters are probably in the very late stages of tidal evolution, which is primarily dominated by $Q^{{}^{\prime}}_{*}$ rather than formation channels. Furthermore, our derived constraint is well consistent with those derived from binary interactions ($\log Q^{{}^{\prime}}_{*}\sim 5.5$; [2005ApJ...620..970M, 2014AJ....148...38M]). In addition, our results emphasizes the necessity of considering the formation pathways and timescales of hot Jupiters when constraining $Q^{{}^{\prime}}_{*}$.
@@
In summary, we find that the declining of the frequency of hot Jupiters with age is broken, which simultaneously puts constraints on the stellar tidal quality factors and the relative importance of the ‘Early’ and ‘Late’ models to form hot Jupiters. Although the secular chaos mechanism [2011ApJ...735..109W, 2017MNRAS.464..688H, 2019MNRAS.486.2265T] is identified as the dominant channel in the ‘Late’ model, the situation regarding the ‘Early’ model (timescales within tenths of Gyr) remains unclear, highlighting the need for a larger sample of hot Jupiters orbiting young stars. Moreover, our results focus primarily on understanding formation channels in single-star systems. Investigating planetary systems in binary star samples can offer further insights into constraining the formation mechanisms (e.g., binary Kozai) of hot Jupiters [e.g., 2024ApJ...977L..11S, 2025ApJ...980L..31W]. Future observations and analyses of hot Jupiters from TESS [2015JATIS...1a4003R], PLATO [2014ExA....38..249R], JWST [2006SSRv..123..485G] and ET 2.0 [2024ChJSS..44..400G] will test our results and further provide more insights on their origin and evolution.